OpenStudy (mysticfury):

What is the direct linear variation equation for the relationship? y varies directly with x and y = 4 when x = 12. A. y = 1/3x B. y = x + 8 C. y = x – 8 D. y = 3x

OpenStudy (photon336):

It's in the form \[y = k*x \]

OpenStudy (anonymous):

A. Solve 4 = k * 12

OpenStudy (mysticfury):

yes, but when i plug in the numbers A and C are both correct

OpenStudy (photon336):

True, but from what I know, direct linear variation is y = k*x

Nnesha (nnesha):

how are you getting A and C

Nnesha (nnesha):

how would you solve for k ?

OpenStudy (mysticfury):

so im going with A then

Nnesha (nnesha):

remember. a/1 isn't same as 1/a

Nnesha (nnesha):

\[\large\rm 4=k *12\] divide both sides by 12 to solve for k(constant of variation ) \[\frac{4}{12}=k\] simplify the fraction

Nnesha (nnesha):

or the other way around as mentioned above equation for direct variation is y=kx substitute in y and x values for both option A and C \[4= \frac{1}{3} \cdot 12 \] \[4=3\cdot 12\] which one is true ?